# Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules

### v2

@@ Acknowledgements.
We thank Adnan Hammud for the Ag-tip fabrication. We thank Jun Yoshinobu and Heiko Appel for helpful discussions. T.S. acknowledges the support of JST FOREST Grant No. JPMJFR221U. T.K. acknowledges the support of JST FOREST Grant No. JPMJFR201J and JSPS KAKENHI Grant No. JP19K24684.References
@@
- Bruch et al. [2007] L. Bruch, R. Diehl, and J. Venables, Progress in the measurement and modeling of physisorbed layers, Rev. Mod. Phys. 79, 1381 (2007).
- Avouris et al. [1982] P. Avouris, D. Schmeisser, and J. Demuth, Observation of rotational excitations of H2 adsorbed on Ag surfaces, Phys. Rev. Lett. 48, 199 (1982).
@@
- Yu et al. [1983] C.-f. Yu, K. B. Whaley, C. S. Hogg, and S. J. Sibener, Selective adsorption resonances in the scattering of -H2, -H2, -D2 and -D2 from Ag(111), Phys. Rev. Lett. 51, 2210 (1983).
@@
- Fukutani et al. [2003] K. Fukutani, K. Yoshida, M. Wilde, W. Diño, M. Matsumoto, and T. Okano, Photostimulated desorption and ortho-para conversion of H2 on Ag surfaces, Phys. Rev. Lett. 90, 096103 (2003).
@@
- Gupta et al. [2005] J. Gupta, C. Lutz, A. Heinrich, and D. Eigler, Strongly coverage-dependent excitations of adsorbed molecular hydrogen, Phys. Rev. B 71, 115416 (2005).
@@
- Thijssen et al. [2006] W. Thijssen, D. Djukic, A. Otte, R. Bremmer, and J. van Ruitenbeek, Vibrationally induced two-level systems in single-molecule junctions, Phys. Rev. Lett. 97, 226806 (2006).
@@
- Trouwborst et al. [2009] M. Trouwborst, E. Huisman, S. van der Molen, and B. van Wees, Bistable hysteresis and resistance switching in hydrogen-gold junctions, Phys. Rev. B 80, 081407(R) (2009).
@@
- Persson and Ryberg [1989] B. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
@@
The experiments were performed in an ultrahigh vacuum (UHV) chamber equipped with an STM-based TERS machine (modified UNISOKU USM-1400) at a sample temperature of 10 K. A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cycles in the chamber. H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was introduced into the STM unit cooled at 10 K via a variable leak valve. The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge, respectively. A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and we used for H2 and for D2 [63]. We used a chemically-etched Ag tip that was further sharpened by focused-ion-beam milling as described elsewhere [40]. The bias voltage was applied to the sample while the tip was grounded. The STM images were obtained in the constant current mode at = 50 pA. The cleanness of the Ag(111) surface and tip was confirmed by STM images [Fig. S1(a)]. We also confirmed that a TERS spectrum of the Ag–Ag junction before hydrogen dosing had no molecule-derived peaks (Fig. S2, at 0 L).
In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was directed through a fused silica window and an Ag-coated parabolic mirror (numerical aperture of 0.6). The incident laser power is 5.9 mW at = 532 nm and 9.7 mW at 633 nm. The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the LT-STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror. The scattered light collected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter, each selected according to the excitation wavelength. For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig. S3), 600 (Figs. 1, 2, and S2), and 1200 line/mm (Fig. 3 and S10). A lower-density grating allows the acquisition of spectra over a wider wavelength range, as shown in Fig. S3 (see Sect. II-B), but resulting in broad and asymmetric Raman peaks. Therefore, for the high-frequency-resolution measurements shown in Fig. 3, the densest grating was used.
@@
DFT calculations were performed with the FHI-aims program package [49, 50]. We used intermediate defaults for basis sets, radial grids and cutoff potential parameters and light defaults for the Hartree multipole expansion and angular grids. This combination proved efficient and reliable against benchmarks to tight settings. We simulated an orthogonal Ag(111) surface cell oriented perpendicular to the axis, containing four layers. The cell parameters used were Å, Å and Å. We employed a -point grid and a dipole correction was employed to maximally decouple periodic images along the direction. We used the PBE exchange-correlation functional and included vdW interactions through the pairwise screened TS-vdWsurf method of Ref. [51] in all simulations reported in the main text. Figure S4 shows a comparison of binding energies with the many-body MBD-NL dispersion correction of Ref. [64], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface. Structures were optimized by allowing only hydrogen molecules and the top layer of Ag atoms of the surface to relax. All the other atoms were kept fixed at their bulk (or tip) positions. When we performed simulations including a model tip structure, this was modeled by an Ag19 cluster at varying distances from the surface. The geometry of the tip is shown in Fig. S5. Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions with molecules close to the tip apex [46, 54, 41].
@@
We inspected binding sites and binding energies of a single adsorbed H2 molecule, finding all values in good agreement with previous studies [52]. The binding energy was defined as the difference between the total energy when the molecule was 10 Å away from the surface and the minimum total energy of the adsorbed molecule. The blue curve in Fig. S4 shows the binding curve of a single H2 molecule adsorbed on the fcc-hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV. The model we used in most simulations contained 11 molecules in the unit cell (see rationale below). For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange curve in Fig. S4) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
@@
Performing vibrational analysis on the optimized single molecule adsorbed on the surface showed that negative frequencies related to rotational and translational modes of the molecule were always present. This led us to consider higher coverages to model local stable clustering of the molecules on the surface. We considered a full monolayer (1 ML), where all fcc-hollow sites of the unit cell were occupied by H2. With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and PIMD at LT ( = 20 K) with the i-PI code [65] connected to FHI-aims. PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and . This number of beads is not fully converged at this temperature even with the colored-noise quantum thermostats, but the purpose of these simulations was exploratory.
@@
While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate. For the quantum case, we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface. We could calculate the radius of gyration of H atoms from the PIMD simulations, belonging to the 3 molecules closest to the tip at any point in the simulation. For the closest tip distance considered ( = 7.4 Å), these were 0.33 Å along and and 0.30 Å along ( = 20 K), indicating a small anisotropy. Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å for H2 and Å for D2 at 10 K. We note that, depending on the density of H2 molecules on the surface, for H2 is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed. These features are disregarded in the current simulations. Further, we note that even though these simulations provide a clear indication that quantum mechanical effects affect the density of the molecular coverage of the surface, a complete determination would require much longer simulations, which would take years to complete or require the development of a tailored machine-learning potential that surpasses the scope of this work.
@@
Figure S6 shows the energy profiles experienced by an H2 molecule when varying the distance of its center of mass with respect to the Ag(111) surface, at different tip-surface distances, and at 0.69 ML coverage. We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium. The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of an H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig. S6). The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions). Figure S7(a) shows a comparison with the limiting case of a single molecule in the unit cell (0.06 ML coverage) at a tip–surface distance of 9.4 Å. In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [17]. As we discuss in the main text, this effect is essential to explain the isotope-dependent red-shift observed experimentally.
@@
Figure S7(b) shows a comparison of the 0.69 ML coverage with the tip fixed at Å and different vdW dispersion corrections, added to the PBE functional. It is clear that details of the vdW dispersion impact the shape of the potential. In this case, the many-body vdW dispersion decreases the distance between the two minima of the asymmetric double-well
@@
where is the potential energy of a molecule with its center of mass fixed at a distance above the substrate surface and with all other coordinates kept at their equilibrium values, as described at the end of Sec. I.2. The vibrational motion is parametrized by the reduced mass and angular frequency . We set the mass of a hydrogen atom to and use .
@@
We computed the numerically exact solutions to the 2D Schrödinger equation corresponding to Eq. (S1) using the sinc-function discrete variable representation (DVR) [66, 67]. The equation was solved for high coverage (0.69 ML) with and . Given the large separation in the timescales of translational and vibrational motion, we also considered the vibrationally adiabatic approximation, whereby one fixes the translational coordinate in Eq. (S1) and solves the corresponding one-dimensional (1D) Schrödinger equation in , yielding the analytical harmonic vibrational energies
These are then used to construct the vibrationally adiabatic translational potential,
@@
We conducted TERS simulations according to the methodology proposed in Refs. [46]. We used Tip A discussed in Ref. [54] and calculated the anharmonic Raman spectrum based on an ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K. We calculated the TERS intensity as , where was obtained by a local polarizability calculation at each geometry in the time-series given by the molecular dynamics. For this part of the calculation, we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface. We note that the method of Ref. [46] is unable to capture the plasmonic enhancement in its totality. We do not observe any noteworthy features in the calculated spectrum and it overall agrees with the experimental spectrum. The red-shift of the rotational band in the simulation is explained by the absence of nuclear spin in the simulations, making calculations reproduce only classical rotations that are red-shifted compared to the quantized transition of para-H2. This simulation is of exploratory character, as achieving statistical convergence through the use of parametrized machine-learned potentials along with accurate plasmonic enhancements is beyond the scope of this paper.
@@
To evaluate the electric field in the STM junction, we performed FEM simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module). We used the three-dimentional model with an Ag tip and Ag sample plate described in Supporting Information of Ref. [40]. To reproduce a plasmonic picocavity in the STM junction, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), an Ag half sphere with a radius of 0.5 nm is attached as an atomic-scale protrusion. For comparison, we also constructed another Ag tip model with an effective radius of 30 nm without the atomic-scale protrusion [see the orange dotted curve in Fig. 2(c) in the main text]. A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 55∘ relative to the sample normal. The perfectly matched layers surrounding the volume are set to absorb all outgoing waves. The dielectric constant of the Ag objects is referenced from the literature [68]. The surface normal component of the electric field is sampled.
An Ag flat plate is placed beneath the tip apex with a gap distance . Here we define the origin of the gap distance ( = 0) as the tip height when the Ag-tip apex forms point contact with the Ag surface. As shown in Table S3, the molecular height is calculated to be 0.32–0.33 nm for D2/Ag(111). In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model. Although the calculated depends on the tip height (Table S3), in the FEM model, we used the constant height for the sampling point at any gap distances. We confirmed that differences in the definition or the molecular position on the order of 0.1 nm do not result in qualitative differences in the field enhancement factor curves. To directly compare with the experimental TERS intensities, the electric field enhancement factor (Ref. [57]) is displayed in the plots of Fig. 2(c) in the main text. To the field enhancement of the nearest neighboring molecule, assuming the molecule is placed in the next adsorption site of Ag(111), = 0.29 nm is used [red triangle in the inset of Fig. 2(c) in the main text]. Thus, when the point at the surface just under the tip is defined as (, , ) = (0, 0, 0), the black and red curves shown in Fig. 2c of the main text represent the plots sampling at (, , ) = (0, 0, 0.20) nm and (0.29, 0, 0.20) nm, respectively.
As shown in Fig. 2, the gap distances with the FEM simulations pm to 850 pm well reproduced the Raman enhancement at the experimental tip height to 0 pm. Considering the Ag-atom radius, for the simulations (Fig. 4 in the main text) is comparable to nm. Therefore, pm corresponds to nm, which was used for the alignment of the experimental and theoretical plots displayed in Fig. 4(e) of the main text.
@@
H2 and D2 have two nuclear spin isomers, ortho and para. para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [69]. para-H2 has only even rotational quantum numbers , whereas and ortho-H2 has only odd . Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4]. In the TERS spectra, we can detect only -H2 (Table S4); as a reference, in the gas phase, ortho-H2 have a Raman-active rotational mode [] at 587 cm-1 [69, 56]. This implies that we detected only the physisorbed species after ortho-para conversion on the surface. In the case of D2, ortho ( = 0, 2) is more stable than para ( = 1) by 7.4 meV and more dominant (67%) in the gas phase at room temperature. Only -D2 was detected in our TERS spectra, indicating again that para-D2 was converted on the surface.
@@
Although we introduced normal hydrogen gases (n-H2 and n-D2) into the chamber, we observed rotational/vibrational transitions only for para-H2 and ortho-D2 in the TERS spectra (Fig. 1 in the main text). The dominance of para-H2 on Ag(111) in the TERS spectra ( = 10 K) is consistent with a previous HREELS study of H2 on an Ag film (exposures of 0.5–10 L, = 10 K) [8], whereas some previous reports detect ortho-H2 on Ag(111). A HREELS study on H2/Ag(111) at 6 K [11] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr). Photostimulated desorption studies [14, 15] shows that at an exposure of 30 L and a temperature of 6 K, ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.

### v3

@@ Acknowledgements.
Acknowledgments—We thank Adnan Hammud for the Ag-tip fabrication. We thank Jun Yoshinobu and Heiko Appel for helpful discussions. T.S. acknowledges the support of JST FOREST Grant No. JPMJFR221U. T.K. acknowledges the support of JST FOREST Grant No. JPMJFR201J.References
@@
- Bruch et al. [2007] L. W. Bruch, R. D. Diehl, and J. A. Venables, Progress in the measurement and modeling of physisorbed layers, Rev. Mod. Phys. 79, 1381 (2007).
- Avouris et al. [1982] P. Avouris, D. Schmeisser, and J. E. Demuth, Observation of rotational excitations of H2 adsorbed on Ag surfaces, Phys. Rev. Lett. 48, 199 (1982).
@@
- Yu et al. [1983] C. F. Yu, K. B. Whaley, C. S. Hogg, and S. J. Sibener, Selective adsorption resonances in the scattering of -H2, -H2, -D2 and -D2 from Ag(111), Phys. Rev. Lett. 51, 2210 (1983).
@@
- Fukutani et al. [2003] K. Fukutani, K. Yoshida, M. Wilde, W. A. Diño, M. Matsumoto, and T. Okano, Photostimulated desorption and ortho-para conversion of H2 on Ag surfaces, Phys. Rev. Lett. 90, 096103 (2003).
@@
- Gupta et al. [2005] J. A. Gupta, C. P. Lutz, A. J. Heinrich, and D. M. Eigler, Strongly coverage-dependent excitations of adsorbed molecular hydrogen, Phys. Rev. B 71, 115416 (2005).
@@
- Thijssen et al. [2006] W. H. A. Thijssen, D. Djukic, A. F. Otte, R. H. Bremmer, and J. M. van Ruitenbeek, Vibrationally induced two-level systems in single-molecule junctions, Phys. Rev. Lett. 97, 226806 (2006).
@@
- Trouwborst et al. [2009] M. L. Trouwborst, E. H. Huisman, S. J. van der Molen, and B. J. van Wees, Bistable hysteresis and resistance switching in hydrogen-gold junctions, Phys. Rev. B 80, 081407(R) (2009).
@@
- Persson and Ryberg [1989] B. N. J. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
@@
The scanning tunneling microscopy (STM) and tip-enhanced Raman spectroscopy (TERS) experiments were performed in an ultrahigh vacuum (UHV) chamber (modified UNISOKU USM-1400) at a sample temperature of 10 K. A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cycles in the chamber. H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was introduced into the STM unit cooled at 10 K via a variable leak valve. The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge, respectively. A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and we used for H2 and for D2 [49]. We used a chemically-etched Ag tip that was further sharpened by focused-ion-beam milling as described elsewhere [40]. The bias voltage was applied to the sample while the tip was grounded. The STM images were obtained in the constant current mode at tunneling current of 50 pA. The cleanness of the Ag(111) surface and tip was confirmed by STM images [Fig. S1(a)]. We also confirmed that a TERS spectrum of the Ag–Ag junction before hydrogen dosing had no molecule-derived peaks (Fig. S2, at 0 L).
In the STM junction, -polarized visible laser (532-nm solid-state or 633-nm HeNe) was directed through a fused silica window and an Ag-coated parabolic mirror. The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the low-temperature (LT) STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror. The scattered light collected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter, each selected according to the excitation wavelength . The nominal incident laser power is 5.9 mW at = 532 nm and 9.7 mW at = 633 nm. For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig. S10), 600 (Figs. 1, 2, and S2), and 1200 line/mm (Fig. 3 and S11). A lower-density grating allows the acquisition of spectra over a wider wavelength range, as shown in Fig. S10 (see Sec. II-B), but resulting in broad and asymmetric Raman peaks. Therefore, for the high-frequency-resolution measurements shown in Fig. 3, the densest grating was used.
@@
Density functional theory (DFT) calculations were performed with the FHI-aims program package [50, 51]. We used intermediate defaults for basis sets, radial grids and cutoff potential parameters and light defaults for the Hartree multipole expansion and angular grids. This combination proved efficient and reliable against benchmarks to tight settings. We simulated an orthogonal Ag(111) surface cell oriented perpendicular to the axis, containing four layers. The cell parameters used were Å, Å and Å. We employed a -point grid and a dipole correction was employed to maximally decouple periodic images along the direction. We used the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and included van der Waals (vdW) interactions through the pairwise screened TS-vdWsurf method of Ref. [52] in all simulations reported in the main text. Figure S3 shows a comparison of binding energies with the many-body dispersion correction termed MBD-NL in Ref. [53], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface. Structures were optimized by allowing only hydrogen molecules and the top layer of Ag atoms of the surface to relax. All the other atoms were kept fixed at their bulk (or tip) positions. When we performed simulations including a model tip structure, this was modeled by an Ag19 cluster at varying distances from the surface. The geometry of the tip is shown in Fig. S4. Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions with molecules close to the tip apex [46, 54, 41].
@@
We inspected binding sites and binding energies of a single adsorbed H2 molecule, finding all values in good agreement with previous studies [55]. The binding energy was defined as the difference between the total energy when the molecule was 10 Å away from the surface and the minimum total energy of the adsorbed molecule. The blue curve in Fig. S3 shows the binding curve of a single H2 molecule adsorbed on the face-centered-cubic (fcc) hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV. The model we used in most simulations contained 11 molecules in the unit cell (see rationale below). For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange curve in Fig. S3) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
@@
Performing vibrational analysis on the optimized single molecule adsorbed on the surface showed that negative frequencies related to rotational and translational modes of the molecule were always present. This led us to consider higher coverages to model local stable clustering of the molecules on the surface. We started by considering a full monolayer (1 ML, i.e., 16 molecules on the surface cell), where all fcc-hollow sites of the unit cell were occupied by H2. With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and path-integral molecular dynamics (PIMD) at LT ( = 20 K) with the i-PI code [56] connected to FHI-aims. PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and . This number of beads is not fully converged at this temperature even with the colored-noise quantum thermostats, but the purpose of these simulations was exploratory.
@@
While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate, leaving only 11 molecules adsorbed, corresponding to a ML coverage. This is the coverage adopted in other simulations in this Letter as the high-density model. For the quantum case, we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface. We could calculate the radius of gyration of H atoms from the PIMD simulations, belonging to the 3 molecules closest to the tip at any point in the simulation. For the closest tip distance considered ( = 7.4 Å), these were 0.33 Å along and and 0.30 Å along ( = 20 K), indicating a small anisotropy. Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å for H2 and Å for D2 at 10 K. We note that, depending on the density of H2 molecules on the surface, for H2 is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed. These features are disregarded in the current simulations. Further, we note that even though these simulations provide a clear indication that quantum mechanical effects affect the density of the molecular coverage of the surface, a complete determination would require much longer simulations, which would take years to complete or require the development of a tailored machine-learning potential that surpasses the scope of this work.
@@
Figure S5 shows the energy profiles experienced by an H2 molecule when varying the distance of its center of mass with respect to the Ag(111) surface, at different tip-surface distances, and at 0.69 ML coverage. We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium. The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of an H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig. S5). The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions). Figure S6(a) shows a comparison with the limiting case of a single molecule in the unit cell (0.06 ML coverage) at a tip–surface distance of 9.4 Å. In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [17]. As we discuss in the main text, this effect is essential to explain the isotope-dependent redshift observed experimentally.
@@
Figure S6(b) shows a comparison of the 0.69 ML coverage with the tip fixed at Å and different vdW dispersion corrections, added to the PBE functional. It is clear that details of the vdW dispersion impact the shape of the potential. In this case, the many-body vdW dispersion decreases the distance between the two minima of the asymmetric double-well. This change would impact the quantitative value of the tip–surface distance at which H2 molecules would prefer to sit closer to the tip, at the low-density limit. We do not expect qualitative changes on the effects we discuss in the main text.
@@
where is the potential energy of a molecule with its center of mass fixed at a distance above the substrate surface and with all other coordinates kept at their equilibrium values, as described at the end of Sec. I.2. The vibrational motion along the intramolecular stretching coordinate is parametrized by the reduced mass and angular frequency . We set the mass of a hydrogen atom to and use , with denoting the mass of the electron.
@@
We computed the numerically exact solutions to the 2D Schrödinger equation corresponding to Eq. (S1) using the sinc-function discrete variable representation (DVR) [57, 58]. The equation was solved for high coverage (0.69 ML) with and . Given the large separation in the timescales of translational and vibrational motion, we also considered the vibrationally adiabatic approximation, whereby one fixes the translational coordinate in Eq. (S1) and solves the corresponding one-dimensional (1D) Schrödinger equation in , with known energy eigenvalues
where is the reduced Planck constant and is the vibrational quantum number. These are then used to construct the vibrationally adiabatic translational potential,
@@
We conducted TERS simulations according to the methodology proposed in Refs. [46]. We used Tip A discussed in Ref. [54] and calculated the anharmonic Raman spectrum based on an ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K. We calculated the TERS intensity as , where was obtained by a local polarizability calculation at each geometry in the time-series given by the molecular dynamics. For this part of the calculation, we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface. We note that the method of Ref. [46] is unable to capture the plasmonic enhancement in its totality. We do not observe any noteworthy features in the calculated spectrum and it overall agrees with the experimental spectrum. The redshift of the rotational band in the simulation is explained by the absence of nuclear spin in the simulations, making calculations reproduce only classical rotations. These are incapable of capturing the quantization of the rotational transitions and thus do not capture the transition of para-H2, where is the rotational quantum number. This explains the redshift observed in the simulated frequency of this transition with respect to the experimental value. This simulation is of exploratory character, as achieving statistical convergence through the use of parametrized machine-learned potentials along with accurate plasmonic enhancements is beyond the scope of this paper.
@@
To evaluate the electric field in the STM junction, we performed finite element method (FEM) simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module). We used the three-dimensional model with an Ag tip and Ag sample plate in a vacuum [40]. The tip shaft is elongated in direction with a half opening angle of 6∘ from a 30-nm radius sphere [Fig. S9(a)]. The tip is axisymmetrical to [see the inset of Fig. 2(c) for the coordinate] with a length of 300 nm. To reproduce a plasmonic picocavity in the STM junction, at the apex of the Ag tip, an Ag half sphere with a radius of 0.5 nm is attached as an atomic-scale protrusion [Fig. S9(b)]. For comparison [the orange dotted curve in Fig. 2(c) of the main text], we also constructed another Ag tip model with an effective radius of 30 nm without the atomic-scale protrusion. A -polarized plane wave with = 532 nm and an electric field of 1 V/m is incident with an angle of 55∘ relative to the axis. The perfectly matched layers surrounding the volume are set to absorb all outgoing waves. The dielectric constant of the Ag objects is referenced from the literature [59]. The component of the electric field is sampled.
An Ag flat plate with a thickness of 100 nm is placed beneath the tip apex with a gap distance , as shown in [Fig. S9(b). As shown in Table S3, the molecular height is calculated to be 0.32–0.33 nm for D2/Ag(111) [Fig. S9(c)]. In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model [the open bullet in Fig. S9(b)]. Although the calculated depends on the tip height (Table S3), the FEM model used the constant height for the sampling point at any . We confirmed that differences in the molecular position on the order of 0.1 nm do not result in qualitative differences in the curve shape. To directly compare with the experimental TERS intensities, the electric field enhancement factor (Ref. [60]) is displayed in Fig. 2(c). To the field enhancement of the nearest neighboring molecule, assuming the molecule is placed in the next adsorption site of Ag(111), = 0.29 nm is used [red triangle in the inset of Fig. 2(c)]. Thus, when the point at the surface just under the tip is defined as (, , ) = (0, 0, 0), the black and red curves shown in Fig. 2c of the main text represent the plots sampling at (, , ) = (0, 0, 0.20) nm and (0.29, 0, 0.20) nm, respectively.
As shown in Fig. 2, the gap distances with the FEM simulations pm to 850 pm well reproduced the Raman enhancement at the experimental tip height to 0 pm. Considering the Ag-atom radius, for the simulations (Fig. 4 in the main text) is comparable to nm [Fig. S9(b) and (c)]. Therefore, pm corresponds to nm, which was used for the alignment of the experimental and theoretical plots of the vibrational frequency shifts [Fig. 4(e)].
@@
H2 and D2 have two nuclear spin isomers, ortho and para. para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [61]. para-H2 has only even , whereas and ortho-H2 has only odd . Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4]. In the TERS spectra, we can detect only -H2 (Table S4); as a reference, in the gas phase, ortho-H2 have a Raman-active rotational mode [] at 587 cm-1 [61, 62]. This implies that we detected only the physisorbed species after ortho-para conversion on the surface. In the case of D2, ortho ( = 0, 2) is more stable than para ( = 1) by 7.4 meV and more dominant (67%) in the gas phase at room temperature. Only -D2 was detected in our TERS spectra, indicating again that para-D2 was converted on the surface.
@@
Although we introduced normal hydrogen gases (n-H2 and n-D2) into the chamber, we observed rotational/vibrational transitions only for para-H2 and ortho-D2 in the TERS spectra (Fig. 1 in the main text). The dominance of para-H2 on Ag(111) in the TERS spectra ( = 10 K) is consistent with a previous study with high-resolution electron energy loss spectroscopy (HREELS) for H2 on an Ag film (exposures of 0.5–10 L, = 10 K) [8], whereas some previous reports detect ortho-H2 on Ag(111). A HREELS study on H2/Ag(111) at 6 K [11] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr). Photostimulated desorption studies [14, 15] shows that at an exposure of 30 L and a temperature of 6 K, ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.